# Automatic BLAS Offloading on Unified Memory Architecture: A Study on NVIDIA Grace-Hopper

### v2

@@ Abstract.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; July 21–25, 2024; Providence, RIisbn: 978-1-4503-XXXX-X/18/06ccs: General and reference Performanceccs: Mathematics of computing Mathematical software performanceccs: Software and its engineering Software performance
@@ 2.1. Coherent memory in NVIDIA Grace-Hopper
In conventional architectures, GPU and host memory exist in separate spaces, preventing direct access between CPU and GPU memory. To alleviate this hurdle, CUDA managed memory introduced since CUDA 6 allows sharing memory pointers, but implicit page migration is performed by the CUDA runtime behind the scene. In contrast, the Grace-Hopper superchip features closely integrated CPU and GPU units, connected by the high-bandwidth and cache-coherent NVLink Chip-2-Chip (C2C) interconnect. This technology enables a unified memory space where both CPU and GPU can access memory without page movement. From the OS perspective, these memories resemble two NUMA domains, akin to memories in a two-socket CPU system.
@@ 4. Performance Testings and Discussion
We focus on exploring the performance on the novel Grace-Hopper (also known as GH200) unified memory architecture. Our Grace-Hopper superchip contains one 72-core Grace CPU with 480GB LPDDR5 host memory and one H100 GPU with 96GB HBM3e memory. Note that the 480GB model has lower bandwidth than models with less LPDDR5, so performance is expected to be better on those models. The GPU driver on GH200 is the latest 550.54.14 and CUDA version is 12.4. Power can be dynamically allocated within the superchip, and TDP of the entire chip is 1000W. The latest NVHPC 24.3 compiler and math libraries in the compiler suite are used throughout the testing.
@@
In all the following test cases, matrix multiplication with problem size $(mnk)^{1/3}>500$ will be offloaded which is proven to be appropriate to benefit from GPU.
@@ 4.1. Matrix Multiplication (dgemm)
Table 3. Runtime of cublasDgemm w/ cudaMemCpy
(transA=’T’, transB=’N’, M=32, N=2400, K=93536)
@@
We tested dgemm call for a matrix size (M=32, N=2400, K=93536) which often appears in the PARSEC application tested in the next subsection. Figure 2 shows performance of using unified memory access with zero data copy (Strategy 2). It’s seen that GPU on HBM is tremendously faster than CPU on LPDDR5, and GPU on LPDDR5 performs similarly to CPU on LPDDR5. However, run CPU on HBM is much slower. This is understandable by the STREAM bandwidth in Table 2, the bandwidth of CPU accessing HBM is quite limited, even slower than accessing LPDDR5. Table 3 shows performance of using explicit cudaMemcpy for every cuBLAS call (Strategy 1), while GH200 and H100-PCIe are both extremely fast in performing dgemm, data transfer in GH200 is much faster than transferring through the PCIe lanes. Note that the cublasDgemm time on cudamalloced memory (0.52ms) in Table 3 is consistently faster than running on host allocated HBM (0.84ms) in Table 2. NVBLAS is also tested on GH200, but significantly slower than our tool. The breakdown timing was measured by internal timer in our tool, so no detailed timing for NVBLAS is available. It is also worth mentioning that CUDA pinned memory is enabled in both our tool and NVBLAS.

### v5

@@ Abstract.
journalyear: 2024copyright: acmlicensedconference: Practice and Experience in Advanced Research Computing; July 21–25, 2024; Providence, RI, USAbooktitle: Practice and Experience in Advanced Research Computing (PEARC ’24), July 21–25, 2024, Providence, RI, USAdoi: 10.1145/3626203.3670561isbn: 979-8-4007-0419-2/24/07ccs: General and reference Performanceccs: Mathematics of computing Mathematical software performanceccs: Software and its engineering Software performance
@@ 2.1. Coherent memory in NVIDIA Grace-Hopper
In conventional architectures, GPU and host memory exist in separate spaces, preventing direct access between CPU and GPU memory. To alleviate this hurdle, CUDA managed memory introduced since CUDA 6 allows sharing memory pointers, but implicit page migration is performed by the CUDA runtime behind the scene. In contrast, the Grace-Hopper superchip features closely integrated CPU and GPU units along with LPDDR5x and HBM3e memory subsystems, connected by the high-bandwidth and cache-coherent NVLink Chip-2-Chip (C2C) interconnect(NVIDIA, 2023). This technology enables a unified memory space where both CPU and GPU can access memory without page movement. From the OS perspective, these memories resemble two NUMA domains, akin to memories in a two-socket CPU system.
@@ 4. Performance Testings and Discussion
We focus on exploring the performance on the novel Grace-Hopper (also known as GH200) unified memory architecture. Our Grace-Hopper superchip contains one 72-core Grace CPU with 480GB LPDDR5x host memory and one H100 GPU with 96GB HBM3e memory. Note that the 480GB model has lower bandwidth than models with less LPDDR5x, so performance is expected to be better on those models. The GPU driver on GH200 is the latest 550.54.14 and CUDA version is 12.4. Power can be dynamically allocated within the superchip, and TDP of the entire chip is 1000W. The latest NVHPC 24.3 compiler and math libraries in the compiler suite are used throughout the testing.
@@
In all the following test cases, matrix multiplication with problem size $(mnk)^{1/3}>500$ will be offloaded which is proven to be appropriate to benefit from GPU. The optimal offload threshold depends on the relative compute capability of GPU versus CPU and latency, and the value can be adjusted by user.
@@ 4.1. Matrix Multiplication (dgemm)
Table 3. Runtime of cublasDgemm w/ cudaMemCpy (transA=’T’, transB=’N’, M=32, N=2400, K=93536)
@@
We tested dgemm call for a matrix size (M=32, N=2400, K=93536) which often appears in the PARSEC application tested in the next subsection. Table 2 shows performance of using unified memory access with zero data copy (Strategy 2). It’s seen that GPU on HBM is tremendously faster than CPU on LPDDR5, and GPU on LPDDR5 performs similarly to CPU on LPDDR5. However, run CPU on HBM is much slower. This is understandable by the STREAM bandwidth in Table 1, the bandwidth of CPU accessing HBM is quite limited, even slower than accessing LPDDR5. Table 3 shows performance of using explicit cudaMemcpy for every cuBLAS call (Strategy 1), while GH200 and H100-PCIe are both extremely fast in performing dgemm, data transfer in GH200 is much faster than transferring through the PCIe lanes. Note that the cublasDgemm time on cudamalloced memory (0.52ms) in Table 3 is consistently faster than running on host allocated HBM (0.84ms) in Table 2. NVBLAS is also tested on GH200, but significantly slower than our tool. The breakdown timing was measured by internal timer in our tool, so no detailed timing for NVBLAS is available. It is also worth mentioning that CUDA pinned memory is enabled in both our tool and NVBLAS.